# Cislunar Constellation Design for Space Situational Awareness with Time-Expanded p-Median Problem

### v2

@@ Abstract
Driven by the surmounting interest for a dedicated infrastructure in cislunar space, this work considers the satellite constellation design for cislunar space situational awareness (CSSA). We propose a linear programming (LP)-based formulation that simultaneously tackles the constellation design and sensor-tasking subproblems surrounding CSSA. Our approach generates constellation designs that provide coverage with considerations for the field-of-view of observers. We propose a time-expanded $p$-Median problem (TE-$p$-MP) which considers the optimal placement of $p$ space-based observers into discretized locations based on orbital slots along libration point orbits, simultaneously with observer pointing directions across discretized time. We further develop a Lagrangian method for the TE-$p$-MP, where a relaxed problem with an analytical solution is derived, and customized heuristics leveraging the orbital structure of candidate observer locations are devised. The performance of the proposed formulation is demonstrated with several case studies for CSSA constellations monitoring the cislunar Cone of Shame and a periodic time-varying transit window for low-energy transfers located in the Earth-Moon L2 neck region. The proposed problem formulation, along with the Lagrangian method, is demonstrated to enable a fast assessment of near-optimal CSSA constellations, equipping decision-makers with a critical technique for exploring the design trade space.
@@ 1.2 Problem Definition and Proposed Approach
While generic BnB algorithms are known to adequately tackle classical $p$-median problems of moderate size, we also propose a custom LM to compute an approximate solution rapidly. LM consists of iteratively solving for an idealized optimal solution to a relaxed problem and a heuristics-based feasible solution, each time updating Lagrange multipliers used to construct the relaxed problem. In this paper, an LM with a purely analytical solution to the relaxed problem is derived by decoupling the observer-wise sensor-tasking decision. Custom heuristics for the feasible solution are designed to leverage the orbital structure of candidate observer locations. We note that the large number of variables, in the order of $10^{6}$, of the TE-$p$-MP prohibits the use of EAs. An NLP-based approach for tackling both constellation design and sensor-tasking problems could consist, for example, of formulating a black-box constellation design problem in the outer loop, with sensor-tasking handled in the inner loop by a pre-trained neural network (NN), or a policy chosen from a predefined set by the outer loop. We however opt for the proposed LP-based approach in favor of its ease in constructing a tractable formulation that encodes decision surrounding the observation schedule. With an NLP-based approach, a policy-based approach would result in a limited performance due to the inherently enforced structure by the pre-defined set of policies, while parameterizing the policy to be optimized by an outer-loop is non-trivial. Recently, Patel et al. [20] also proposed a framework for concurrent CSSA constellation design and sensor-tasking. Their approach is a hybrid between NLP and LP, where an outer-layer NLP optimizes the phasing of satellites along pre-defined orbits, and an inner-layer LP that assigns discrete targets to observers. This work differs from [20] in the following ways: our formulation optimizes not only for the phasing but also the orbits of observers from a discrete set of options; we also consider target observation by considering a sensor FOV, rather than assuming a one-to-one assignment between observers and targets; finally, our formulation is a single-layer MILP that does not rely on a black-box NLP solver. Through the proposed TE-$p$-MP together with the LM, our approach can produce a feasible candidate CSSA constellation design together with a pointing direction schedule for each observer that provides adequate coverage to a given set of demand and observer parameters.
@@ 2.1.1 $p$-Median Problem
where $f_{j}$ is the deployment cost of a candidate facility location $j$, and $c_{ij}$ denote the allocation cost of a facility at $j$ to client $i$. The objective (3a) is based on the cumulative sum of the deployment costs of the $p$ facilities and the total allocation cost; constraints (3b) ensure each client is serviced by exactly one facility; constraint (3c) ensures exactly $p$ facility locations are used; constraints (3d) ensure if at least one client is allocated to facility location $j$, then there is a facility at $j$; finally, constraints (3e) ensure $X_{ij}$ and $Y_{j}$ are binary. The TE-$p$-MP formulated in this work is based on (3), but with appropriate modifications made to (1) accommodate an objective function based on cumulative target coverage and to (2) include time indices.
@@ 2.3.2 Libration Point Orbit
This work considers the synodic-resonant LPOs in the L1/L2 Lyapunov, Southern and Northern L2 Halo, Southern and Northern Butterfly, Distant Prograde Orbit (DPO), and Distant Retrograde Orbit (DRO) families, with synodic resonances 1:1, 3:2, 2:1, 9:4, 5:2, 3:1, 4:1, and 9:2. This selection results in 40 LPOs, as shown in Figure 1. The parameters of these LPOs are given in Table 7 in Appendix 7.1. Each LPO is constructed via a fixed-time variant of the single-shooting differential correction approach [58], using the LPO entry in the JPL Three-Body Periodic Orbit Catalog [^4] https://ssd.jpl.nasa.gov/tools/periodic_orbits.html with the closest period to the sought resonant period as initial guess. The choice of which resonant LPOs to include as candidate orbits matters if the relative positions of member satellites within a constellation must be repeated over a $N_{\rm syn}$-multiple of the synodic month, where $N_{\rm syn}$ is the smallest integer such that the product $N_{\rm syn}M/N$ is an integer for all considered $M$:$N$ ratios. This translates to whether the sought observer location and pointing schedule are only required to be transient, or if a repeating, steady-state solution is sought. With our choice of resonant LPOs, the constellation returns exactly to its initial configuration every 4 synodic months. We note that the TE-$p$-MP formulation developed in later Sections does not explicitly depend on the $N_{\rm syn}$-month repeating pattern of candidate LPOs, and thus one may instead choose to use an arbitrary set of candidate observer LPOs. However, the number of discretized time steps necessitated to solve for a CSSA constellation and sensor-tasking over an extended duration, for example, lasting multiple years, can be reduced by several folds.
@@
Figure 1: Synodic-resonant LPOs considered as candidate observer orbits. For the Halo and Butterfly families, only the Southern branch is shown, but the Northern branch is also included as candidate orbits.
@@ 3 Demand and Visibility Definitions
We introduce the demands considered for the CSSA problem. Specifically, we first distinguish between static and dynamic demands; the former consists of coverage requirement that spans the same volume over time, while the latter consists of time-dependent coverage, either to keep custody of known targets or to monitor transfer corridors that are time-dependent. In both cases, demands are defined by considering a finite set of targets distributed in space; in the static case, the demand for all targets is assumed to be always active, while in the dynamic case, the demand is activated only at time steps when the target exists. Consider $k=1,\ldots,q$ targets and $t=1,\ldots,\ell$ time steps; we define the demand matrix $D\in\mathbb{B}^{\ell\times q}$ as
@@
Figure 2: Cislunar observation targets for persistent monitoring, shown in Earth-Moon rotating frame centered at the Earth-Moon barycenter. Red shaded volume represents the region of interest approximated by the discrete set of targets.
@@ 3.1 Static Demand: Cone of Shame
A static demand is defined by a set of targets distributed in space that remain stationary. While in reality, an object cannot remain at a fixed location in space, each “target” in this scenario may be understood as placeholder for defining a volume of interest. Then, the coverage of such a static demand translates to perpetual coverage of the volume that is spanned by the spatial distribution of these static targets.
@@ 3.2 Dynamic Demand: Low-Energy Transfer Transit Window
Figure 3: Subset of 1000 sample LETs from pre-computed database in the Earth-Moon rotating frame centered at the Earth-Moon system barycenter. Grey markers indicate transit window monitoring targets.
Figure 4: Target sparsity pattern of demand matrix $D\in\mathbb{B}^{\ell\times q}$ for persistent coverage of low-energy transfer arrival window over a span of 120 time steps lasting for four synodic months
@@ 3.3.1 Sun Illumination Model
In this work, we consider targets with $R_{\rm targ}=2\,\mathrm{m}$, $C_{\rm diff}=0.2$, and $C_{\rm spec}=0$. Figure 5 shows the contour of $\bar{m}_{\mathrm{target}}$ centered at the observer with $\phi_{S}=200^{\circ}$; the contour demonstrates the strong variation of $\bar{m}_{\mathrm{target}}$ with respect to $\phi_{S}$, and the weaker variation with respect to the distance $\|\boldsymbol{r}_{k}-\boldsymbol{r}_{j}\|$ as the latter increases.
Figure 5: Example visibility magnitude contour with $\phi_{S}=200^{\circ}$, with Sun illumination direction aligned with orange arrow.
@@ 3.3.3 Visibility boolean
Figure 7: Illustration of definitions used to compute the visibility boolean of a target $k$ from an observer location $j$
@@ 4.2 Problem Formulation
The objective (22a) has two terms; the primary objective is the first term, which maximizes the fraction of visible targets at each time-step, and takes its summation. Note that (22a) is different from simply taking $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}$, which would allow the optimizer to capitalize on observing the same target by multiple observers rather than observing as many targets as possible. The second term is a discount factor that penalizes the use of “expensive” observer locations, where $f\in\mathbb{R}^{n}$ is the vector of costs associated with making use of a given facility location, given by
@@
where $\nu_{j}$ is the stability index of the LPO, as defined by equation (14). The definition of $f_{j}$ with equation (23) allows the “cost” of a given facility location to have values close to one another: for the considered set of LPOs, $\nu_{j}$ varies between $1$ and $1399.19$, while $f_{j}$ varies between about $0.9091$ and $0.9993$. Since prohibitively expensive candidate observer locations can be excluded from the candidate observer location set $\mathcal{J}$ to begin with, it is assumed that all candidate LPOs have acceptable station-keeping cost should they be chosen as part of the constellation. As such, the second term in the objective is simply meant to favor combinations of facilities that are more stable in the case that multiple combinations of facilities yield the same value from the first term.
Constraint (22b) ensures each observer location $j$ only looks to up to a single direction $i$ at any given time $t$. Constraint (22c) is the traditional $p$-median constraint, which ensures exactly $p$ observer is placed. Constraint (22d) is also a common constraint in the traditional $p$-median problem, enforcing the $j^{\mathrm{th}}$ observer to exist for its allocation to view direction $i$ at time $t$ to be used in the solution. Finally, constraint (22e) ensures the variables $X$ and $Y$ are binary.
@@
where constraint (24b) ensures this modified problem (24) to have the same optimal solution as the original problem (22). Problem (24) is a standard mixed-binary linear program (MBLP), which is amenable to standard solvers such as Gurobi [65] or CPLEX [66]. While the slack variable $\theta_{tk}$ is a continuous variable in $[0,1]$, the optimal solution to (24) has $\theta_{tk}=1$ for time $t$ and target $k$ if at least one observer is viewing $k$ at this time, and $\theta_{tk}=0$ if $\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}=0$. Thus, this slack variable may be interpreted as a binary variable, such that
@@ 5.1 Upper Bound Solution via Lagrangian Relaxation
We first seek to find the decision variables and corresponding upper bound solution to the relaxed problem. Relaxing constraints (22b) and (24b) with multipliers $\lambda_{jt}\geq 0$ and $\eta_{tk}\geq 0$ respectively, the relaxed problem is given by
@@ 5.1.2 Construction of Lagrangian Relaxed Solution
$Z^{\pi(1)}_{\mathrm{relax}}\geq Z^{\pi(2)}_{\mathrm{relax}}\geq\ldots\geq Z^{\pi(m)}_{\mathrm{relax}}$ (33)
@@ 5.3 Heuristics for Neighborhood Facility Swap
where $\texttt{Neighborhood}(j)$ is a set of slot indices that are in the “neighborhood” of slot $j$. We conceive two kinds of “neighborhoods”, one within the same LPO as $j$ but with shifted phasing, and another with similar phasing but on different LPOs. These are defined in further details in the subsequent subsections.
@@ 5.3.1 Intra-Orbit Neighborhood
The intra-orbit neighborhood is defined based on proximity in angular location along a given LPO. Specifically, let $\Xi_{j}\subset\mathcal{J}$ denote all other facility indices that are on the same LPO as $j$; for a given $j$, its LPO is discretized into $b_{j}$ slots according to equation (13), then $\Xi_{j}$ has $b_{j}-1$ elements, excluding $j$ itself. We further define $\Delta\varphi_{j\xi}$ as the angular separation between facility locations $j$ and $\xi\in\Xi_{j}$. Then, we define the permutation of $\xi$, denoted by $\pi_{\rm intra}$, for which the absolute value of $\Delta\phi_{j\xi}$ is in ascending order,

### v3

# Cislunar Space Situational Awareness Constellation Design and Planning with Facility Location Problem

@@ Abstract
Driven by the surmounting interest for dedicated infrastructure in cislunar space, this work considers the satellite constellation design for cislunar space situational awareness (CSSA). We propose a mixed-integer linear programming (MILP)-based formulation that simultaneously tackles the constellation design and sensor-tasking subproblems surrounding CSSA. Our approach generates constellation designs that provide coverage with considerations for the field-of-view of observers. We propose a time-expanded $p$-Median problem (TE-$p$-MP) which considers the optimal placement of $p$ space-based observers into discretized locations based on orbital slots along libration point orbits, simultaneously with observer pointing directions across discretized time. We further develop a Lagrangian method for the TE-$p$-MP, where a relaxed problem with an analytical solution is derived, and customized heuristics leveraging the orbital structure of candidate observer locations are devised. The performance of the proposed formulation is demonstrated with several case studies for CSSA constellations monitoring the cislunar Cone of Shame and a periodic time-varying transit window for low-energy transfers located in the Earth-Moon L2 neck region. The proposed problem formulation, along with the Lagrangian method, is demonstrated to enable a fast assessment of near-optimal CSSA constellations, equipping decision-makers with a critical technique for exploring the design trade space.
@@ 1.2 Problem Definition and Proposed Approach
While generic BnB algorithms are known to adequately tackle classical $p$-median problems of moderate size, we also propose a custom LM to compute an approximate solution rapidly. LM consists of iteratively solving for an idealized optimal solution to a relaxed problem and a heuristics-based feasible solution, each time updating Lagrange multipliers used to construct the relaxed problem. In this paper, an LM with a purely analytical solution to the relaxed problem is derived by decoupling the observer-wise sensor-tasking decision. Custom heuristics for the feasible solution are designed to leverage the orbital structure of candidate observer locations. We note that the large number of variables, in the order of $10^{6}$, of the TE-$p$-MP prohibits the use of EAs. An NLP-based approach for tackling both constellation design and sensor-tasking problems could consist, for example, of formulating a black-box constellation design problem in the outer loop, with sensor-tasking handled in the inner loop by a pre-trained neural network (NN), or a policy chosen from a predefined set by the outer loop. We however opt for the proposed LP-based approach in favor of its ease in constructing a tractable formulation that encodes decisions surrounding the observation schedule. With an NLP-based approach, a policy-based approach would result in a limited performance due to the inherently enforced structure by the pre-defined set of policies, while parameterizing the policy to be optimized by an outer loop is non-trivial. Recently, Patel et al. [20] also proposed a framework for concurrent CSSA constellation design and sensor-tasking. Their approach is a hybrid between NLP and LP, where an outer-layer NLP optimizes the phasing of satellites along pre-defined orbits, and an inner-layer LP assigns discrete targets to observers. This work differs from [20] in the following ways: our formulation optimizes not only for the phasing but also the orbits of observers from a discrete set of options; we also consider target observation by considering a sensor FOV, rather than assuming a one-to-one assignment between observers and targets; finally, our formulation is a single-layer MILP that does not rely on a black-box NLP solver. Through the proposed TE-$p$-MP together with the LM, our approach can produce a feasible candidate CSSA constellation design together with a pointing direction schedule for each observer that provides adequate coverage to a given set of demand and observer parameters.
@@ 2.1.1 $p$-Median Problem
where $f_{j}$ is the deployment cost of a candidate facility location $j$, and $c_{ij}$ denote the allocation cost of a facility at $j$ to client $i$. The objective (3a) is based on the cumulative sum of the deployment costs of the $p$ facilities and the total allocation cost; constraints (3b) ensure each client is serviced by exactly one facility; constraint (3c) ensures exactly $p$ facility locations are used; constraints (3d) ensure if at least one client is allocated to facility location $j$, then there is a facility at $j$; finally, constraints (3e) ensure $X_{ij}$ and $Y_{j}$ are binary. The TE-$p$-MP formulated in this work is based on (3), but with appropriate modifications made (1) to accommodate an objective function based on cumulative target coverage and (2) to include time indices. The use of an equality constraint in (3c) instead of an inequality constraint of the form $\sum_{j=1}^{n}Y_{j}\leq p$ is reasonable for problem scenarios where adding an additional facility can always improve the objective. The equality form will be particularly useful when deriving the Lagrangian relaxation in Section 5.1.
@@ 2.3.2 Libration Point Orbit
This work considers the synodic-resonant LPOs in the L1/L2 Lyapunov, Southern and Northern L2 Halo, Southern and Northern Butterfly, Distant Prograde Orbit (DPO), and Distant Retrograde Orbit (DRO) families, with synodic resonances 1:1, 3:2, 2:1, 9:4, 5:2, 3:1, 4:1, and 9:2. This selection results in 40 LPOs, as shown in Figure 1. The parameters of these LPOs are given in Appendix 7.1. Each LPO is constructed via a fixed-time variant of the single-shooting differential correction approach [58], using the LPO entry in the JPL Three-Body Periodic Orbit Catalog [^4] https://ssd.jpl.nasa.gov/tools/periodic_orbits.html with the closest period to the sought resonant period as initial guess. The choice of which resonant LPOs to include as candidate orbits matters if the relative positions of member satellites within a constellation must be repeated over a $N_{\rm syn}$-multiple of the synodic month, where $N_{\rm syn}$ is the smallest integer such that the product $N_{\rm syn}M/N$ is an integer for all considered $M$:$N$ ratios. This translates to whether the sought observer location and pointing schedule are only required to be transient, or if a repeating, steady-state solution is sought. With our choice of resonant LPOs, the constellation returns exactly to its initial configuration every 4 synodic months. We note that the TE-$p$-MP formulation developed in later Sections does not explicitly depend on the $N_{\rm syn}$-month repeating pattern of candidate LPOs, and thus one may instead choose to use an arbitrary set of candidate observer LPOs. However, the number of discretized time steps necessitated to solve a CSSA constellation and sensor-tasking over an extended duration, for example, lasting multiple years, can be reduced by several folds.
@@
Figure 1: Synodic-resonant LPOs considered as candidate observer orbits, showing only Southern branch for Northern/Southern symmetric LPOs
@@ 3 Demand and Visibility Definitions
We introduce the demands considered for the CSSA problem. Specifically, we first distinguish between static and dynamic demands; the former consists of coverage requirement that spans the same volume over time, while the latter consists of time-dependent coverage, either to keep custody of known targets or to monitor transfer corridors that are time-dependent. In both cases, demands are defined by considering a finite set of targets distributed in space, which approximates a volume of interest for monitoring. In the static case, the demand for all targets is assumed to be always active, while in the dynamic case, the demand is activated only at time steps when the target exists. Consider $k=1,\ldots,q$ targets and $t=1,\ldots,\ell$ time steps; we define the demand matrix $D\in\mathbb{B}^{\ell\times q}$ as
@@
Figure 2: Cislunar observation targets for persistent monitoring of volume of interest (red shaded volume), shown in Earth-Moon rotating frame centered at the Earth-Moon barycenter
@@ 3.1 Static Demand: Cone of Shame
A static demand is defined by a set of targets distributed in space that remain stationary. While in reality, an object cannot remain at a fixed location in space, each “target” in this scenario may be understood as a placeholder for defining a volume of interest. Then, the coverage of such a static demand translates to perpetual coverage of the volume that is spanned by the spatial distribution of these static targets.
@@ 3.2 Dynamic Demand: Low-Energy Transfer Transit Window
Figure 3: Sample LETs in the Earth-Moon rotating frame centered at the system barycenter. Grey markers indicate transit window monitoring targets.
Figure 4: Sparsity of demand matrix $D\in\mathbb{B}^{\ell\times q}$ for LET transit window over 120 time steps corresponding to $4$ synodic months
@@ 3.3.1 Sun Illumination Model
In this work, we consider targets of $R_{\rm targ}=2\,\mathrm{m}$, approximating a medium-sized, robotic lunar spacecraft, with $C_{\rm diff}=0.2$, and $C_{\rm spec}=0$. Values for $C_{\rm diff}$ and $C_{\rm spec}$ are according to [2], and correspond to a Lambertian sphere model. An object may exhibit a higher $C_{\rm diff}$, for example within the range $[0.1,1]$ as stated in [6], and the selected value serves as a conservative estimate, which results in dimmer targets. Figure 5 shows the contour of $\bar{m}_{\mathrm{target}}$ centered at the observer with $\phi_{S}=200^{\circ}$; the contour demonstrates the strong variation of $\bar{m}_{\mathrm{target}}$ with respect to $\phi_{S}$, and the weaker variation with respect to the distance $\|\boldsymbol{r}_{k}-\boldsymbol{r}_{j}\|$ as the latter increases.
Figure 5: Example visibility magnitude contour with $\phi_{S}=200^{\circ}$, with Sun illumination direction aligned with orange arrow
@@ 3.3.3 Visibility boolean
Figure 7: Discretizations used to compute the visibility boolean of a target $k$ from an observer location $j$
@@ 4.2 Problem Formulation
The objective (22a) has two terms; the primary objective is the first term, which corresponds to the number of observed targets across all time steps. Note that the first term in (22a) is different from simply taking $\sum_{t=1}^{\ell}\sum_{k=1}^{q}\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}$, which would allow a solution to capitalize on observing the same target by multiple observers rather than observing as many targets as possible. The second term is a discount factor that penalizes the use of “expensive” observer locations in the case that multiple combinations of facilities yield the same value from the first term. The vector $f\in\mathbb{R}^{n}$, defined by
@@
models the cost of a given facility location $j$, where $\nu_{j}$ is the stability index of the LPO, as defined by equation (14). The definition of $f_{j}$ with equation (23) allows the “cost” of a given facility location to be distributed around similar values, all under $1$: for the considered set of LPOs, $\nu_{j}$ varies between $1$ and $1399.19$, while $f_{j}$ varies between about $0.9091$ and $0.9993$. Directly mapping $f_{j}=\nu_{j}$ would be inappropriate for the objective (22a), as the “cost” associated with placing a single observer on an unstable LPO, say of $\nu_{j}=1399.19$, would be approximately the same as placing $1399$ observers on a stable LPO of $\nu_{j^{\prime}}=1$, which is clearly unrealistic. In effect, by defining $f_{j}\in[0,1)$, we ensure the unstable LPOs, some which are known be advantageous for CSSA [2, 13], are not discouraged if a better target coverage can be achieved. Since prohibitively expensive candidate observer locations can be excluded from the candidate observer location set $\mathcal{J}$ to begin with, it is assumed that all candidate LPOs have acceptable station-keeping cost should they be chosen as part of the constellation. To ensure the optimal solution never compromises on the number of observed targets to choose a “cheaper” constellation according to costs $f_{j}$, the second term is multiplied by $1/\ell$. Since $f_{j}<1\,\forall j$, $\sum_{j=1}^{n}f_{j}Y_{j}<p$ due to (22c). Also, for a typical CSSA problem, $p<<\ell$, so $(1/\ell)\sum_{j=1}^{n}f_{j}Y_{j}<p/\ell<<1$. Thus, an additional target observation at any single time-step is always prioritized in an optimal solution to (22).
Constraint (22b) ensures each observer location $j$ only looks to up to a single direction $i$ at any given time $t$. Constraint (22c) is the traditional $p$-median constraint, which ensures exactly $p$ observer is placed. As mentioned in Section 2.1.1, the use of an exact equality constraint in (22c) is due to the assumption that for cislunar SSA where the coverage demand is high, having an additional observer can always improve the objective (22a). Constraint (22d) is also a common constraint in the traditional $p$-median problem, enforcing the $j^{\mathrm{th}}$ observer to exist for its allocation to view direction $i$ at time $t$ to be used in the solution. Finally, constraint (22e) ensures the variables $X$ and $Y$ are binary.
@@
where constraint (24b) ensures this modified problem (24) has the same optimal solution as the original problem (22). Problem (24) is a standard mixed-integer linear program (MILP), which is amenable to standard BnB solvers such as Gurobi [65] or CPLEX [66]. While the slack variable $\theta_{tk}$ is a continuous variable in $[0,1]$, the optimal solution to (24) has $\theta_{tk}=1$ for time $t$ and target $k$ if at least one observer is viewing $k$ at this time, and $\theta_{tk}=0$ if $\sum_{i=1}^{m}\sum_{j=1}^{n}M^{b}_{ijtk}X_{ijt}=0$. Thus, this slack variable may be interpreted as a binary variable, such that
@@ 5.1 Upper Bound Solution via Lagrangian Relaxation
We first seek to find the decision variables and the corresponding upper-bound solution to the relaxed problem. Relaxing constraints (22b) and (24b) with multipliers $\lambda_{jt}\geq 0$ and $\eta_{tk}\geq 0$ respectively, the relaxed problem is given by
@@ 5.1.2 Construction of Lagrangian Relaxed Solution
$Z^{\pi(1)}_{\mathrm{relax}}\geq Z^{\pi(2)}_{\mathrm{relax}}\geq\ldots\geq Z^{\pi(n)}_{\mathrm{relax}}$ (33)
@@ 5.3 Heuristics for Neighborhood Facility Swap
where $\texttt{Neighborhood}(j)$ is a set of slot indices that are in the “neighborhood” of slot $j$. We conceive two kinds of “neighborhoods”, one within the same LPO as $j$ but with shifted phasing, and another with similar phasing but on different LPOs. These are defined in further detail in the subsequent subsections.
@@ 5.3.1 Intra-Orbit Neighborhood
The intra-orbit neighborhood is defined based on proximity in angular location along a given LPO, as shown in Figure 9(a). Specifically, let $\Xi_{j}\subset\mathcal{J}$ denote all other facility indices that are on the same LPO as $j$; for a given $j$, its LPO is discretized into $b_{j}$ slots according to equation (13), then $\Xi_{j}$ has $b_{j}-1$ elements, excluding $j$ itself. We further define $\Delta\varphi_{j\xi}$ as the angular separation between facility locations $j$ and $\xi\in\Xi_{j}$. Then, we define the permutation of $\xi$, denoted by $\pi_{\rm intra}$, for which the absolute value of $\Delta\phi_{j\xi}$ is in ascending order,